Opened 7 years ago

Last modified 5 years ago

#377 assigned flight processing

UR10/01 flight day 259b/2010, Iceland

Reported by: anch Owned by: knpa
Priority: alpha 5 Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: /data/okhotsk1/scratch/arsf/archive-overflow/UR10_01-2010_259b_Iceland

Data arrived from ARSF via SATA disk DAPHNE 03 on 7/10/2010

Scientific objective: Not known

Priority: A5

PI: M. Kirkbride

Sensors:

  • Eagle (25/02/11)
  • Hawk (25/02/11)
  • Leica LIDAR (11/01/11)
  • Digital Photography (11/01/11)

Change History (41)

comment:1 Changed 7 years ago by anch

  • Description modified (diff)

comment:2 Changed 7 years ago by iopa

Applanix and IPAS done and are in 259a in the workspace.

comment:3 Changed 7 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Starting eagle processing.

comment:4 Changed 7 years ago by knpa

Found eagle scts for most lines but a couple don't have sufficient wobbles.

Need LIDAR to correct.

comment:5 Changed 7 years ago by knpa

Taking LIDAR.

comment:6 Changed 7 years ago by knpa

Processed LIDAR has gaps in the data over the high areas, presumably volcano. Could be due to smoke or ice but looks to me like a range issue. Need to investigate further. The range settings in alspp look fine but Mark1 says there is a separate 'range gate' that is set manually. A record of this is kept in flightlinelog in leica/logs but this seems to be missing. Will request this file from ops and investigate further.

Note, this also occured in the other sortie of the day, 259a.

comment:7 Changed 7 years ago by anch

This has also occurred on flight day 258 (Am updating the other ticket now).

comment:8 Changed 7 years ago by knpa

Phone call from Phil:

This is indeed an issue with the range gate, the plane was flying lower than intended relative to the ground due to an out of date DTM (the volcano had caused the terrain to rise).

Can't do anything about these gaps, resuming processing.

comment:9 Changed 7 years ago by iopa

Camera delivery has been checked.

It might be worth adding a sentence in data quality remarks about images 622-625,629-631,
705-712 saying that the images are correct but the high brightness reflected from the snow has
caused the lens' shape to appear in the recorded image.

Otherwise all is good and ready for delivery.

comment:10 Changed 7 years ago by knpa

LIDAR QC:

Vertical and horizontally discrepancies OK. Doesn't appear to be any significant consistent roll.

Currently classifying. There is a LOT of noise on all flightlines (due to Ice I think, this would be consistent with previous projects).

comment:11 Changed 7 years ago by knpa

Finished LIDAR - awaiting delivery check.

comment:12 Changed 7 years ago by anch

Delivery checking LIDAR

comment:13 Changed 6 years ago by anch

LIDAR delivery checked - all looks good to go except the file

dem/README_DEM_MAY_NOT_BE_ACCURATE.txt

Which mentions SRTM when ASTER data has been used to generate the DEM.

comment:14 Changed 6 years ago by knpa

Have fixed above, preparing lidar for delivery.

comment:15 Changed 6 years ago by knpa

LIDAR and Photography data sent to M. Kirkbride at Dundee on 11/01/11

comment:16 Changed 6 years ago by knpa

  • Description modified (diff)

comment:17 Changed 6 years ago by anch

Starting Hawk processing

comment:18 Changed 6 years ago by anch

Submitted for delivery check

comment:19 Changed 6 years ago by knpa

Eagle SCTS:

1 -0.01
2 0.00*
3 0.02
4 -0.035*
5 0.00
6 0.03
7 -0.01*
8 -0.01
9 0.03
10 0.00
11 0.00
12 0.03
13 0.04
14 0.03
15 0.01
16 0.01
17 -0.01*
18 0.00
19 0.01
20 0.00*
21 0.00
22 -0.02*
23 0.03
24 -0.01*
  • indicates this figure was found by matching with the lidar. The rest were checked for distortions within wobbles.

comment:20 Changed 6 years ago by knpa

Submitted Eagle for delivery check.

comment:21 Changed 6 years ago by anch

Merged Hawk and Eagle deliveries into single hyperspectral delivery.

comment:22 Changed 6 years ago by iopa

Delivery checking Hyperspectral.

comment:23 Changed 6 years ago by iopa

Hyperspectral delivery check;

  • the logsheet calls the flight day '259a page 2' but every other part of the project correctly uses 259b
  • the geoid-spheroid should be contained in the misc/ when there is no LiDAR dem

Otherwise the delivery seems to be in order and ready to go.

comment:24 Changed 6 years ago by knpa

There was a couple of issues still with the Readme which anch has fixed.

Preparing HS for delivery, will go with 248 Hawk which is also ready.

comment:25 Changed 6 years ago by knpa

  • Description modified (diff)

Eagle and Hawk delivered to M. Kirkbride at Dundee (25/02/11)

comment:26 Changed 5 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Milestone set to 2010 data processing completion

Preparing for archiving.

comment:27 Changed 5 years ago by knpa

This project still needs to be rsynced which will temporarily use up more space on okhotsk. Will leave this one until we have a bit more breathing space.

comment:28 Changed 5 years ago by knpa

Appears to be r-synced now (not sure who by), so proceeding with archiving.

comment:29 Changed 5 years ago by knpa

Hawk SCTs (from .sync file)

HAWK

1 0.00
2 0.00
3 0.00
4 0.03
5 0.00
6 0.00
7 0.00
8 0.01
9 0.00
10 0.00
11 0.01
12 0.02
13 0.00
14 0.00
15 0.00
16 0.00
17 0.00
18 0.00
19 0.00
20 0.00
21 -0.03
22 0.00
23 0.00
24 0.00

comment:30 Changed 5 years ago by knpa

navigation intermediate files are missing.

comment:31 Changed 5 years ago by knpa

There is some FWF data for this project.

There is tabi data for this project.

comment:32 Changed 5 years ago by knpa

The lidar delivery screenshots are missing.

comment:33 Changed 5 years ago by knpa

Found the above screenshots - they were in the HS delivery for some bizarre reason.

comment:34 Changed 5 years ago by knpa

Removed eagle delivery 20110204 because as far as I can tell this was never delivered and was actually just merged with the main HS delivery 20110211.

comment:35 Changed 5 years ago by knpa

Make available to NEODC 24/07/12

comment:36 Changed 5 years ago by knpa

Delivery checking reprocessed hyperspectral by jaho.

comment:37 Changed 5 years ago by knpa

Delivery check complete.

Many issues (incl some major) will need to be fixed by someone before delivering:

  1. No lev3 mapped file for e12, e16, e18 e19
  2. Need to rename screenshots so they match the lev1 files 1-24 (instead of the raw files 11-34)
  3. Following screenshots missing: e29 h23
  4. These screenshots show a large gap in the middle (probably linked to the lack of mapped files above): e26, e28
  5. The following files have file size discrepancies: e1, e2 (mask), e3 (nav), e4 (nav), e10, e12(nav), e13, e14, e14 (mask), h1, h1 (mask bad pixel), h2 (mask), h3 (nav), h10 (mask bad pixel), h13 (# lines in igm does not agree with # in lev1)
  6. Typo in ReadMe data quality section "flighlines". Should also specify that they are REGULAR dropped scan lines.
  7. apltran example command has an incorrect filename format for the output file
  8. The logsheet filename is incorrect
  9. SCT error on hawk 06

I will try and correct all these seeing as jaho has left.

comment:38 Changed 5 years ago by adbe

Hyperspectral reprocessing re-delivery check: No problems, ready to deliver

comment:39 Changed 5 years ago by knpa

PI has not requested a re-delivery so this is finished with.

comment:40 Changed 5 years ago by knpa

Original deliveries archived by NEODC.

Moved to archive disk.

comment:41 Changed 5 years ago by knpa

Removed workspace version (original processing).

Leaving ticket open until reprocessing is archived.

Note: See TracTickets for help on using tickets.